# When can in-context learning generalize out of task distribution?

### v1

@@ 1 Introduction
Prior work (Raventos et al., 2023; Lu et al., 2024; He et al., 2024) has focused on understanding how the number of pretraining tasks affects the ability of the model to generalize to unseen tasks (generated from the same distribution as the pretraining tasks). Here, we ask a related but distinct question: If a model is pretrained only on tasks from a subset of the full task space, what conditions are necessary for it to generalize to the rest of the space? We think of this question as asking about the out-of-distribution (OOD) performance of models trained to do in-context learning. This question prompts us to consider a more general notion of task diversity that depends not only on task enumeration but also on how different or similar they are. Sampling a distribution with many similar tasks has the potential to induce the model towards a more specialized ICL solution that performs well only on novel tasks within its pretraining distribution. However, we observe that transformers trained to do ICL of linear functions undergo a transition from a specialized solution to one that generalizes over the full task space as we increase the degree of task diversity. This phenomenon of out-of-distribution task generalization sheds new light on in-context learning behavior.
@@ 1.3 Contributions
  • We examine how two notions of task diversity (number of tasks and task similarity) interact, and construct a phase diagram that reveals three distinct regimes of ICL generalization.
@@ 3 Experimental Results
Unless stated otherwise, we study $d=10$ dimensional regression with $n=50$ examples in each context. We use a GPT-2 style transformer (Radford et al., 2019) with learned positional embeddings, a hidden dimension of $d_{h}=128$, 10 layers (except in Fig 9), and 8 attention heads. We use a learned linear embedding to map $x_{i}$ and $y_{i}$ to the hidden dimension $d_{h}=128$. The target values $y_{i}$ are padded with $d-1$ zeroes. For further training details, see Appendix section A.1.
@@ 3.2 Specialization-generalization transition
Figure 6: Phase diagrams reveal three phases of generalization. A: Phase diagram for in-task-distribution test loss ($\delta=0^{\circ}$). B: Phase diagram for out-of-task-distribution test loss ($\delta=175^{\circ})$. Diagonal structure reveals interplay between two the measures of task diversity (Training $\phi$, number of tasks $N$). C: Combining the earlier two phase diagrams reveals three phases of in-context learning. In-weights learning, teal: the model fits the training data but fails to generalize, both in- and out-of-task-distribution. In-task-distribution generalization, yellow: The model generalizes within the support of the pretraining distribution, but fails to generalize out of distribution. Out-of-task-distribution generalization, purple: The model generalizes well both in- and out-of-task-distribution. In constructing the phase diagram, we set the threshold between high and low generalization losses to $10^{-2}$.
@@ 4 Discussion and future work
[…]
@@ 4.1 Types of distribution shift
This is not the only type of distribution shift one may care about, however, and prior work has investigated ICL under different distribution shifts. Ahuja & Lopez-Paz (2023) investigate instead a domain shift of the $x_{i}$’s, (i.e. $p(x_{i})\rightarrow p^{\prime}(x_{i})$ and find that transformers fail to generalize when exposed to such shifts. We suggest that this may be because the authors do not expose the model to sufficient data diversity during pretraining. In Appendix A.6 we investigate the effect of data diversity in the linear regression setting and see a (noisy) transition in the behavior of the model’s OOD generalization. Yadlowsky et al. (2024, 2023) instead consider ICL under concept shift: the mapping of data to labels $p(y_{n}|C_{n})$ changes between train and test. The authors here find that transformers largely fail to generalize to such distribution shifts, even when exposed to a diversity of tasks during pretraining. The ability our models display in successfully adapting to distribution shift between train and test time may therefore depend on the type of distribution shift considered, in addition to task/data diversity thresholds.
@@ A.5 Sampling from a portion of the hypersphere
WLOG, we take $v=R\hat{e}_{1}=(R\ \vec{0})$, where $\vec{0}\in\mathbb{R}^{d}$. Then, writing $w=(Z\ y)$ with $Z\in\mathbb{R}$ and $y\in\mathbb{R}^{d}$, we see that
@@
We will show that, up to this restriction, $Z=R(2U-1)$, with $U\sim\operatorname{Beta}(\frac{d}{2},\frac{d}{2})$. To generate $y$, we first draw $Z$, then draw a vector uniformly from the sphere $S^{d-1}(\sqrt{R^{2}-Z^{2}})$
@@
$Z=R(2U-1)$, with $U\sim\operatorname{Beta}(\frac{d}{2},\frac{d}{2})$
@@
Making the change of variables $Z=R(2u-1)$, we obtain:

### v2

@@ 1 Introduction
Prior work (Raventos et al., 2023; Lu et al., 2024; He et al., 2024) has focused on understanding how the number of pretraining tasks affects the ability of the model to generalize to unseen tasks (generated from the same distribution as the pretraining tasks). Here, we ask a related but distinct question: If a model is pretrained only on tasks from a subset of the full task space, what conditions are necessary for it to generalize to the rest of the space? We think of this question as asking about the out-of-distribution (OOD) performance of models trained to do in-context learning. This question prompts us to consider a more general notion of task diversity that depends not only on task enumeration but also on how different or similar they are. Sampling a distribution with many similar tasks has the potential to induce the model towards a more specialized ICL solution that performs well only on novel tasks within its pretraining distribution. However, we observe that transformers trained to do ICL of linear functions undergo a transition from a specialized solution to one that generalizes over the full task space as we increase the degree of task diversity. This phenomenon of out-of-distribution task generalization sheds new light on in-context learning behavior [^1] Code available at https://github.com/cwgoddard/OOD_ICL.
@@ 1.3 Contributions
  • We examine how two notions of task diversity (number of tasks and task similarity [^3] For linear problems, it is natural to choose the inner product between tasks $w_{1}^{T}w_{2}$ as a similarity measure.) interact, and construct a phase diagram that reveals three distinct regimes of ICL generalization.
@@ 3 Experimental Results
Unless stated otherwise, we study $d=10$ dimensional regression with $n=50$ examples in each context. We use a GPT-2 style transformer (Radford et al., 2019) with learned positional embeddings, a hidden dimension of $d_{h}=128$, 10 layers (except in Fig 9), and 8 attention heads. We use a learned linear embedding to map $x_{i}$ and $y_{i}$ to the hidden dimension $d_{h}=128$. The target values $y_{i}$ are padded with d1d-1 zeroes. For further training details, see Appendix section A.1.
@@ 3.2 Specialization-generalization transition
Figure 6: Phase diagrams reveal three phases of generalization. A: Phase diagram for in-task-distribution test loss ($\delta=0^{\circ}$). B: Phase diagram for out-of-task-distribution test loss ($\delta=175^{\circ})$. Diagonal structure reveals interplay between two the measures of task diversity (Training $\phi$, number of tasks NN). C: Combining the earlier two phase diagrams reveals three phases of in-context learning. In-weights learning, teal: the model fits the training data but fails to generalize, both in- and out-of-task-distribution. In-task-distribution generalization, yellow: The model generalizes within the support of the pretraining distribution, but fails to generalize out of distribution. Out-of-task-distribution generalization, purple: The model generalizes well both in- and out-of-task-distribution. In constructing the phase diagram, we set the threshold between high and low generalization losses to $10^{-2}$.
@@ 4 Discussion and future work
Figure 10: Task distribution diversity induces a transition from specialized to general-purpose ICL in nonlinear regression tasks. A: All parameters in the nonlinear model (a small one-hidden-layer network) are drawn from the same hypersphere. The transition occurs at $\phi\approx 60^{\circ}$. Although the curves for each $\delta$ do not collapse onto each other after the transition, they remain within a band of size $\sim$$10^{-3}$. B: The parameters in the nonlinear model are drawn separately from a different hypersphere for each layer in the model. The transition occurs at $\phi\approx 135^{\circ}$.
@@ 4.1 Types of distribution shift
This is not the only type of distribution shift one may care about, however, and prior work has investigated ICL under different distribution shifts. Ahuja & Lopez-Paz (2023) investigate instead a domain shift of the $x_{i}$’s, (i.e. $p(x_{i})\rightarrow p^{\prime}(x_{i})$ and find that transformers fail to generalize when exposed to such shifts. We suggest that this may be because the authors do not expose the model to sufficient data diversity during pretraining. In Appendix A.6 we investigate the effect of data diversity in the linear regression setting and see a (noisy) transition in the behavior of the model’s OOD generalization. Yadlowsky et al. (2024, 2023) instead consider ICL under concept shift: the mapping of data to labels $p(y_{n}|C_{n})$ changes between train and test. The authors here find that transformers largely fail to generalize to such distribution shifts, even when exposed to a diversity of tasks during pretraining. Hill et al. (2025) consider ICL of linear functions where task vectors at test time are orthogonal to those seen during training, and again see that transformers fail to generalize to this distribution shift. We expect that this is because the natural task similarity measure for linear tasks, the inner product, is zero for orthogonal tasks. The ability our models display in successfully adapting to distribution shift between train and test time may therefore depend on the type of distribution shift considered, in addition to task/data diversity thresholds.
@@ A.5 Sampling from a portion of the hypersphere
WLOG, we take $v=R\hat{e}_{1}=(R\ \vec{0})$, where $\vec{0}\in\mathbb{R}^{d}$. Then, writing w=(Zy)w=(Z\ y) with $Z\in\mathbb{R}$ and $y\in\mathbb{R}^{d}$, we see that
@@
We will show that, up to this restriction, Z=R(2U1)Z=R(2U-1), with $U\sim\operatorname{Beta}(\frac{d}{2},\frac{d}{2})$. To generate $y$, we first draw $Z$, then draw a vector uniformly from the sphere $S^{d-1}(\sqrt{R^{2}-Z^{2}})$
@@
Z=R(2U1)Z=R(2U-1), with $U\sim\operatorname{Beta}(\frac{d}{2},\frac{d}{2})$
@@
Making the change of variables Z=R(2u1)Z=R(2u-1), we obtain: